(a) Give two examples of continuous time discrete state random processes.
(b) Find the mean, autocorrelation and autocovariance of the random process defined by X(t) =
A cos ωt, where A is uniformly distributed in (0,1) and ω is a constant.
1
Expert's answer
2021-08-16T09:44:07-0400

Part a)

Discrete-time continuous state Markov processes are widely used. Autoregressive processes are a very important example.

Actually, if you relax the Markov property and look at discrete-time continuous state stochastic processes in general, then this is the topic of study of a huge part of Time series analysis and signal processing.

The most famous examples are ARMA processes, the Conditionally Heteroscedastic models, a large subclass of Hidden Markov models


Part b)

Given, random process, "X(t) = Kcos(wt) t \u2265 0"

The expected value of K is, "E[K]=\frac{1+0}{2}=0.5"

The variance of K is, "\sigma_K^2 = \frac{(1-0)^2}{12}=\frac{1}{12}"

Thus

"E[K^2] = \sigma_K^2+ [E[K]]^2 \\
E[K^2] = \frac{1}{12} + (0.5)^2 \\
E[K^2] = \frac{1}{ 3}"

The mean of X(t) is,

"E[X(t)] = E[Kcos(wt)] \\
E[X(t)] = E[K]cos(wt) \\
E[X(t)] = \frac{1}{2} cos(wt)"


The autocorrelation function of X(t) is,

"R_{xx}(t, s) = E[X(t)X(s)] \\
R_{xx}(t, s) = E[K^2cos(wt)cos(ws)] \\
R_{xx}(t, s) = E[K^2]cos(wt)cos(ws)\\
 R_{xx}(t, s) = \frac{1}{3} cos(wt)cos(ws)"


The autocovariance function of X(t) is,

"C_{xx}(t, s) = R_{xx}(t, s) — E[X(t)]E[X (s)]\\
C_{xx}(t, s) = \frac{1}{3}cos(wt)cos(ws) -\frac{1}{4} cos(wt)cos(ws)\\
 C_{xx}(t, s) = \frac{1}{12}cos(wt)cos(ws)"